Gallium has two naturally occurring isotopes with mass 69 and 71 respectively. What is the peecentage abundance of both isotopes if their relative atomic mass is 70.59?
1
Expert's answer
2021-08-01T19:37:02-0400

solution

Let the percentage of the isotope 71 Gallium be =y%

The percentage of the isotope 69Ga is =(100−y) %

Therefore, we write the mass balance equation

100×69.72 = y×71+(100−y) ×69

6972 = 71y+6900−69y

2y=6972−6900=72

y=36


The percentage of the isotope is =36%